# MonoLSS: Learnable Sample Selection For Monocular 3D Detection

### v1

@@ Abstract
In the field of autonomous driving, monocular 3D detection is a critical task which estimates 3D properties (depth, dimension, and orientation) of objects in a single RGB image. Previous works have used features in a heuristic way to learn 3D properties, without considering that inappropriate features could have adverse effects. In this paper, sample selection is introduced that only suitable samples should be trained to regress the 3D properties. To select samples adaptively, we propose a Learnable Sample Selection (LSS) module, which is based on Gumbel-Softmax and a relative-distance sample divider. The LSS module works under a warm-up strategy leading to an improvement in training stability. Additionally, since the LSS module dedicated to 3D property sample selection relies on object-level features, we further develop a data augmentation method named MixUp3D to enrich 3D property samples which conforms to imaging principles without introducing ambiguity. As two orthogonal methods, the LSS module and MixUp3D can be utilized independently or in conjunction. Sufficient experiments have shown that their combined use can lead to synergistic effects, yielding improvements that transcend the mere sum of their individual applications. Leveraging the LSS module and the MixUp3D, without any extra data, our method named MonoLSS ranks 1st in all three categories (Car, Cyclist, and Pedestrian) on KITTI 3D object detection benchmark, and achieves competitive results on both the Waymo dataset and KITTI-nuScenes cross-dataset evaluation. The code is included in the supplementary material and will be released to facilitate related academic and industrial studies.
@@ 1 Introduction
  • To enrich 3D property samples, we devise MixUp3D data augmentation, which simulates spatial overlap and significantly improves 3D detection performance.
@@ 4.3 Ablation Study
Effectiveness of the LSS Module. As reported in the first 4 rows of Table 5, the LSS module can significantly improve the AP. While the improvement is relatively tiny when it acts on the orientation and dimension property. This is because the depth estimation error is the most critical limiting factor in monocular 3D detection, which has been identified in GeoAug [29] and MonoDLE [36]. Thus, for the convenience of comparison, the LSS module only acts on the depth property in the following ablation experiments.
@@
Synergistic effect between LSS and MixUp3D. As presented in Table 5, while the LSS module and MixUp3D each exhibit a positive impact when applied independently ($+3.06$ and $+3.8$ in Easy level), their combined usage results in a greater improvement ($+11.74$) than the sum of their individual effects ($+6.68$), indicating a mutual enhancement effect between them.

### v2

@@ Abstract
In the field of autonomous driving, monocular 3D detection is a critical task which estimates 3D properties (depth, dimension, and orientation) of objects in a single RGB image. Previous works have used features in a heuristic way to learn 3D properties, without considering that inappropriate features could have adverse effects. In this paper, sample selection is introduced that only suitable samples should be trained to regress the 3D properties. To select samples adaptively, we propose a Learnable Sample Selection (LSS) module, which is based on Gumbel-Softmax and a relative-distance sample divider. The LSS module works under a warm-up strategy leading to an improvement in training stability. Additionally, since the LSS module dedicated to 3D property sample selection relies on object-level features, we further develop a data augmentation method named MixUp3D to enrich 3D property samples which conforms to imaging principles without introducing ambiguity. As two orthogonal methods, the LSS module and MixUp3D can be utilized independently or in conjunction. Leveraging the LSS module and the MixUp3D, without any extra data, our method named MonoLSS ranks 1st in all three categories (Car, Cyclist, and Pedestrian) on KITTI 3D object detection benchmark, and achieves competitive results on both the Waymo dataset and KITTI-nuScenes cross-dataset evaluation. The code is available at https://github.com/Traffic-X/MonoLSS.
@@ 1 Introduction
  • To enrich 3D property samples, we devise MixUp3D data augmentation, which simulates spatial overlap and improves 3D detection performance.
@@ 4.3 Ablation Study
Effectiveness of the LSS Module. As reported in the first 4 rows of Table 5, the LSS module can significantly improve the AP. While the effect is slightly off when it acts on the orientation and dimension property. This is because the depth estimation error is the most critical limiting factor in monocular 3D detection, which has been identified in GeoAug [29] and MonoDLE [36]. Thus, for the convenience of comparison, the LSS module only acts on the depth property in the following ablation experiments.
@@
Effect between LSS and MixUp3D. As presented in Table 5, while the LSS module and MixUp3D each exhibit a positive impact when applied independently ($+3.06$ and $+2.93$ in Easy level), their combined usage results in a higher improvement ($+4.19$).